an isosceles triangle has two sides with a measure of 16 inches. The base of the triangle is 12 inches. Find the height of the triangle.
Answer by Boreal(15235)

The sides are equal, so the height is over the middle of the base.
This splits the original triangle into two, with side 6 (half the base) and hypotenuse 16.
Pythagorean theorem: 6^2+x^2=16^2
36+x^2=256
x^2=220
x=sqrt (220)=2 sqrt (55) inches